A string is streched b\w 2 fixd pts separated by 75 cm!it is observed to have resonant frequencies of 420Hz and 315Hz!~there are no resonant freq between these two. then the fundamental frequency for the string is______________!!1!!!pls ans !!!i'l approve the answer if it's right!!!!!!!!!

Grade:12

1 Answers

harini desiraju
13 years ago

420=n+1/2(l)[V]  .......(1)

315=n/2(l)[V]  ........(2)

(1)/(2)

n/n+1=3/4. Considering least possible case,n=3.

420=4(100)/2(75)[V]

V=210

so,fundamental frequency=100/2(75)[210]=140Hz